an arrow is fired with at an angle of 50 degrees with an initial velocity of 300 m/s. It travelled 900 meters before hitting the ground. How long is the bullet in the air?


1
Expert's answer
2021-03-08T08:27:55-0500
"x=v_0tcos\theta,""t=\dfrac{x}{v_0cos\theta}=\dfrac{900\ m}{300\ \dfrac{m}{s}\cdot cos50^{\circ}}=4.67\ s."